Accelerator Physics (T. Satogata)
Accelerator R and D (Y. Zhang)
Computational Physics (Y. Roblin)
- Received a new version of GHOST. It is multiGPU and can do gear changing spend the last couple of weeks trying to understand it and make comparisons with Hao's paper.
- Writing code to output the relevant physics quantities.
- Compiling a list of features that are needed for the next release
- Assisted Fanglei on setting up ELEGANT calculations on the qcd farm
- Cooler meetings
- Reviewing responses from gmn collaboration regarding readiness review.
- Metings on feedback systems and preparing for the ops retreat
- Compiling information and having discussions on the talks for JPOS regarding
- The use of CEBAF as a positron machine
- Reviewed paper for Phys. Rev Letter.
Next two weeks:
- meeting for JPOS (coordination of speakers for the various talks)
- Continuing work on GHOST
- Getting ready for new hire.
- Further work on PERLE lattice
- Working on user interface/data input for JSPEC. Added a parser to be able to specify certain quantities in terms of variables and formulas.
- Wrote a makefile to make it easy to build the code under linux.
- DC cooling at 8GeV: summarized results, put them into slides.
- Summer student Zachary finished implemented IBS calculation formulas. About to start the final step of the project which is the numerical integration. He is also working on the abstract for his report.
- Reviewed paper for Journal of Computational Physics.
Next two weeks -
- Work with Zachary to help him finish his summer project
- Continue working on user interface for JSPEC
- Work on linac optimization paper
- Start/continue work on modifying model beam model to include effects of smaller e- beam on the ion beam its an item in the FY17 supplemental )
Diagnostic Development (K. Jordan)
- Sorting through 20 years of stuff in the FEL trailers in preparation of the August eviction
- Completed work on two CRADAs; ~$170k RadiaBeam wire scanner & ~$80k AlphaCore rad hard high speed camera. This will now take 6 - 8 weeks to drag it’s way through the halls of bureaucracy…
- ‘Nearly’ completed new isotope production rig (image below), & worked with Hari to prepare proposal for Arne.
- Had discussions with Mike & Joe on sync light & pin hole X-Ray cameras.
- Discussed with Brian the possibility of using gated ProSilica cameras for GTS data acquisition
- ISR3C12 - resolved issues with how to move forward in the future - it was installed without mirror as per Arne
I*SR2C20 - resolved issues with installation & alignment
- GTS Viewers
- GTS fiber collection optics (more to follow in another email), Joe’s ray trace design was >20x the collection over Pavel’s design - this calls into question the work that was done at HZDR on wire scanner OTR
- GTS Harps
- New Isotope Production assembly model. FEA to follow
- Worked with George K. on importing Fluka load distributions. Worked well after I wrote a translation algorithm.
- Fixed the 3D printer, realigned the bed
- Made an optical centering rig for conflat flanges from 6” to 2-3/4”
- I have been spending my time on diagnostic related items the past few weeks. I mentioned to Kevin the (to me interesting) observations I'd been making on synchrotron radiation.
- Mostly I've been thinking about and preparing for:
- StayTreat talk on "Unbalancing the Linacs" (coming up rapidly in early Aug)
- JPos17 talk on CEBAF magnet issues with reversed polarity (coming up but not quite as rapidly for Sept)
- Diagnostic options for SR (REALLY fascinating)
- a beam line port for fairly hard S.R. X-rays (10's of keV)
- pinhole photography
- real, refractive focusing
- using off-the-shelf CCD cameras for direct X-ray imaging
- We may be able to do harp-based X-ray focus detection, scanning the harp wire along the focal plane of a refractive X-ray lens. It may be instructive to compare the profile from a wire with different elemental composition than an attenuator plate against the profile from a wire of the same material as the plate. The X-ray spectrum should have holes in it at the atomic inner-shell absorption energies, and these holes may be useful to do spectrally specific profiles and avoid some chromatic effects in the lenses.
- I've also been collaborating with Joe on diagnostic layout issues and interpretation of potential past errors. We'd like not to repeat any errors, and even avoid new ones....
LERF (S. Benson)
July 10-July 21
- Hosted two July Cooler ring meeting (July 12 and 19). Wrote up minutes.
- Finally got ANL undulator jaws sent to ANL.
- Certified User Lab 4 for the HIPPO laser.
- Received documentation from BNL on cooler solenoid specifications.
- Completed the Drive laser LOSP and submitted it for review.
- Helped set up the agenda for the Neil symposium.
- Prepared input for the DoE quarterly report.
- Worked on organizing sensitive, FOUO, ITAR, or export controlled information in Trailer 41C for eventual storage in the Arc building.
- Spent a lot of time going over equipment in Trailer 41B and 41C, especially Michelle Shinn's records. Made good progress but there is still a lot to clean up.
July 24 - August 4th
- Finish packing up office and start move to Arc building.
- Continue to clear out Trailer 41B.
- Host Cooler ring meetings on July 26th and August 2nd and document
- Try to understand cooler solenoid specifications from Brookhaven.
- Finish preparing and help present talk at the Stay-treat on uses of the LERF for CEBAF operations and reliability.
- Attend the Stay-treat, August 2-3.